# Creativity in AI: Progresses and Challenges

### v2

@@ 3.2. Figurative Language
3.3. Lexical Innovation
Understanding and generating novel words or word compounds is a challenging linguistic task that often requires creativity, commonsense knowledge, and an ability to generalize over seen concepts (Costello and Keane, 2000; Wisniewski, 1997). Similar noun compounds might have different meanings based on our common understanding. For example, knowing that “chocolate croissant” means a “croissant filled with chocolate” does not necessarily imply that “chocolate bunny” would mean “a bunny filled with chocolate”, but rather a bar of “chocolate in the shape of a bunny”. Several works have evaluated and analyzed language models on the task of interpreting and predicting the emergence of these noun compounds and found that models generally show a moderate performance (Coil and Shwartz, 2023; Dhar and van der Plas, 2019; Kuznetsova et al., 2013). Other works have successfully trained neural networks to generate neologisms (i.e. newly coined words or phrases) (Das and Ghosh, 2017; Lencione et al., 2022). On the other hand, previous works have also shown that large language models can fail at linguistic generalization tasks such as morphologically deriving new words from nonce roots (Ismayilzada et al., 2024; Weissweiler et al., 2023) and can occasionally duplicate large amounts of text from its training data (McCoy et al., 2021). Similarly, a recent work explores the linguistic creativity of both large language models and humans by reconstructing their text output from the existing text snippets on the web and finds that the seemingly remarkable creativity of model outputs may be in large part attributable to the remarkable creativity of human-written texts on the web (Lu et al., 2024b).
4. Creative Problem-Solving
@@ 4.1. Convergent Thinking
Language models have recently been evaluated on problems that require convergent thinking. Lin et al. (2021) tests language models on solving riddles that require creativity and commonsense and finds a significant gap between model and human performance. Naeini et al. (2023) uses the popular British quiz show Only Connect’s Connecting Wall that mimics RAT formulation with built-in, deliberate red herrings (i.e. misleading stimuli or distractors) and evaluates large language models such as GPT-4 on these problems. They report poor model performance and show that models are highly susceptible to distractors in the input and manifest a form of fixation effect (a.k.a functional fixedness or Einstellung effect). This type of cognitive bias forces the model to fixate on its past knowledge and prevents it from thinking “out-of-the-box”. The same effect is also found when models are evaluated on everyday problems involving unconventional use of objects (Tian et al., 2023). Very recently, large language models such as GPT-4o have been evaluated on the popular New York Times game Connections and have been found to struggle with associating encyclopedic and linguistic knowledge at an abstract level (Samadarshi et al., 2024). Another study has revealed that language models also fall short of demonstrating human-like creativity in code generation (Lu et al., 2024c).
4.2. Divergent Thinking
@@
While language models perform strongly on AUT-like divergent thinking tasks, they, however, struggle when these tasks require some form of lateral thinking or “thinking out-of-the-box”. For example, recent works have found that defying default commonsense associations and modeling unexpected or unlikely situations are challenging for large language models (Jiang et al., 2023; Tian et al., 2023; Zhao et al., 2023a). Figure 4(a) illustrates a creative problem-solving example from Tian et al. (2023) that involves unconventional use of everyday objects.
4.3. Abstraction and Analogy-Making
@@ 4.3.1. Abstraction
As the cornerstone of human intelligence, abstraction, and abstract reasoning are typically evaluated using visual IQ tests in humans. Popular examples of these tests are RAVEN progressive matrices (Raven, 1938), Bongard problems (Bongard, 1970) and recently introduced Kandinsky Patterns (Holzinger et al., 2019), Abstraction and Reasoning Corpus (ARC) (Chollet, 2019) and its variations (Moskvichev et al., 2023). These tests require the participants to identify and complete an abstract visual pattern based on given examples. Although several attempts have been made to solve these tasks using both symbolic-based and neural network-driven approaches (Lorello et al., 2024; Mirchandani et al., 2023; Hu et al., 2023; Xu et al., 2022; Santoro et al., 2018), modern AI systems still struggle with solving RAVEN-like (Ahrabian et al., 2024; Gendron et al., 2023; Odouard and Mitchell, 2022; Zhang et al., 2019a) and ARC-like tasks (Moskvichev et al., 2023; Mitchell et al., 2023; Xu et al., 2023; Odouard and Mitchell, 2022; Zhang et al., 2021). Analysis of abstraction via a serial reproduction task (Langlois et al., 2021) where participants are asked to produce a textual stimulus for the next participant upon observing a visual stimulus and vice versa, has suggested that GPT-4 unlike humans relies heavily on linguistic representations even in vision-only paradigm (Kumar et al., 2024). Figure 5 illustrates an example from the ARC task (Chollet, 2019). The problems in this corpus are quite hard to solve to the extent that this task has been recognized as the de facto benchmark for measuring progress towards Artificial General Intelligence (AGI) and a public competition with a grand prize of $ has recently been launched222https://arcprize.org/. At the time of writing this paper, the highest score is far from from the passing threshold of (human-level).
4.3.2. Analogy-Making
@@
In its basic form, analogy-making is the ability to identify a relation between two concepts and apply it to a new concept. For example, Paris is to France as Tokyo is to Japan (i.e. capital:country relation). Early approaches to computational analogy-making were symbolic-based and required extensive hand-coded input i.e. structured representations of both the entities and their relations (Turney, 2008; Gentner, 1983; Falkenhainer et al., 1989). Later, word embedding models based on neural networks were shown to exhibit analogy-making abilities at the word level (Jacob et al., 2023; Petersen and van der Plas, 2023; Marquer et al., 2022; Ushio et al., 2021; Gladkova et al., 2016; Mikolov et al., 2013). However, word-level analogies are generally simple and do not reflect the typical analogical reasoning humans perform in everyday life about complex situations. While some works have argued for emergent analogical reasoning abilities of large language models (Webb et al., 2022; Yasunaga et al., 2023; Hu et al., 2023), other works have shown that these models lack the robustness and generality exhibited by humans when it comes to long text analogies (Wijesiriwardene et al., 2023; Zhu and de Melo, 2020), scientific analogies (Yuan et al., 2023; Czinczoll et al., 2022), story analogies (Jiayang et al., 2023; Sourati et al., 2023; Sultan and Shahaf, 2022; Nagarajah et al., 2022), visual analogies (Opielka et al., 2024; Bitton et al., 2022; Zhang et al., 2019a) and complex analogical reasoning (Musker et al., 2024; Lewis and Mitchell, 2024). Figure 4(b) illustrates an example from the analogical reasoning over narratives benchmark (Sourati et al., 2023).
5. Artistic Creativity
@@ 5.1. Story Generation
Language models have also been evaluated on their ability to produce and judge creative content as a professional writer (Chakrabarty et al., 2023b; Marco et al., 2024). Chakrabarty et al. (2023b) generates short stories from LLMs based on the plots of popular fictional stories published in the New York Times and conducts a fine-grained expert assessment of both model-generated and original stories. Their study shows that LLMs significantly lag behind seasoned writers in producing inherently creative content. Studies also demonstrate that LLMs are unreliable evaluators of creativity (Chhun et al., 2024; Chakrabarty et al., 2023b). Additionally, (Tian et al., 2024) finds that LLM-generated stories are positively homogenous and typically lack suspense and tension.
@@ 6. Scientific Creativity
Another line of work has focused on automating the discovery of other scientific knowledge such as generating new mathematical conjectures or theories (Raayoni et al., 2021; Wu and Tegmark, 2019; Chen et al., 2016; Buchberger et al., 2006; Fajtlowicz, 1988), automatically proving theorems (Hubert et al., 2024b; Trinh et al., 2024), discovering new concepts (Hakuk and Reich, 2020; Iten et al., 2020; Lenat and Brown, 1984) and predicting molecular structures (Abramson et al., 2024; Zambaldi et al., 2024; Jumper et al., 2021; Lindsay, 1980) among others. A notable example in this area is the recent AlphaFold model (Jumper et al., 2021) that can predict millions of intricate 3D protein structures which has the potential to significantly accelerate research in biology.
@@
While the latest advancements in the automation of scientific creativity are remarkable, these results should be taken with a grain of salt. Most of the recent end-to-end automation frameworks are powered by LLMs, hence, they face the same challenges and issues we discussed in the previous sections such as hallucinations, lack of content diversity, novelty, and robust reasoning capabilities. For example, one of the aforementioned large-scale idea generation studies (Si et al., 2024) finds that out of LLM-generated ideas only are unique. Their qualitative analysis also reveals some common failure modes such as vague implementation details, misuse of datasets, inappropriate baselines, unrealistic assumptions, and overall poorly-motivated ideas. Similarly, another study benchmarking the machine learning experimentation capabilities of LLMs reports hallucinations and poor planning as some of the major issues with these models (Huang et al., 2024b).
@@ 8.1.2. Dimensions of Creativity
As we discussed earlier, there are many dimensions of creativity, but most works generally focus on evaluation of the novelty and usefulness dimensions. However, surprise, agency and spontaneity dimensions are also equally important. Humans typically communicate an emotion or a deeper meaning through creative products and their creative process is characterized by spontaneous “Aha” or “Eureka” moments coupled with deliberate decisions made at each step of the way 666https://www.newyorker.com/culture/the-weekend-essay/why-ai-isnt-going-to-make-art. However, current AI systems lack agency and are typically trained to generate the most likely output leaving no room for any intentional or spontaneous action (Franceschelli and Musolesi, 2023). Therefore, a holistic evaluation of machine creativity should involve consideration of all these different dimensions that characterize human creativity.
@@ Acknowledgements.
We gratefully acknowledge the support of Swiss National Science Foundation (grant 205121_207437: C - LING).References
@@
- Arnheim (2001) Rudolf Arnheim. 2001. What it Means to be Creative. British Journal of Aesthetics 41 (01 2001). https://doi.org/10.1093/bjaesthetics/41.1.24
@@
- Barron (1955) Frank Barron. 1955. The disposition toward originality. The Journal of Abnormal and Social Psychology 51, 3 (1955), 478.
@@
- Elgammal et al. (2017) Ahmed Elgammal, Bingchen Liu, Mohamed Elhoseiny, and Marian Mazzone. 2017. CAN: Creative Adversarial Networks, Generating ”Art” by Learning About Styles and Deviating from Style Norms. arXiv:1706.07068 [cs.AI]
@@
- Huang et al. (2024b) Qian Huang, Jian Vora, Percy Liang, and Jure Leskovec. 2024b. MLAgentBench: Evaluating Language Agents on Machine Learning Experimentation. arXiv:2310.03302 [cs.LG] https://arxiv.org/abs/2310.03302
@@
- Ismayilzada et al. (2024) Mete Ismayilzada, Defne Circi, Jonne Sälevä, Hale Sirin, Abdullatif Köksal, Bhuwan Dhingra, Antoine Bosselut, Lonneke van der Plas, and Duygu Ataman. 2024. Evaluating Morphological Compositional Generalization in Large Language Models. arXiv:2410.12656 [cs.CL] https://arxiv.org/abs/2410.12656
@@
- Stein (1953) Morris I Stein. 1953. Creativity and culture. The journal of psychology 36, 2 (1953), 311–322.
@@
- Weisberg (2015a) Robert Weisberg. 2015a. On the Usefulness of “Value” in the Definition of Creativity. Creativity Research Journal 27 (06 2015), 111–124. https://doi.org/10.1080/10400419.2015.1030320
- Weisberg (2015b) Robert W. Weisberg. 2015b. On the Usefulness of “Value” in the Definition of Creativity. Creativity Research Journal 27, 2 (2015), 111–124. https://doi.org/10.1080/10400419.2015.1030320 arXiv:https://doi.org/10.1080/10400419.2015.1030320
@@
- Xie et al. (2023b) Binzhu Xie, Sicheng Zhang, Zitang Zhou, Bo Li, Yuanhan Zhang, Jack Hessel, Jingkang Yang, and Ziwei Liu. 2023b. FunQA: Towards Surprising Video Comprehension. ArXiv abs/2306.14899 (2023). https://api.semanticscholar.org/CorpusID:259262297
@@
- Xie et al. (2023a) Zhuohan Xie, Miao Li, Trevor Cohn, and Jey Lau. 2023a. DeltaScore: Fine-Grained Story Evaluation with Perturbations. In Findings of the Association for Computational Linguistics: EMNLP 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 5317–5331. https://doi.org/10.18653/v1/2023.findings-emnlp.353

### v3

@@ 3.1.2. Figurative Language
3.1.3. Lexical Innovation
Understanding and generating novel words or word compounds is a challenging linguistic task that often requires creativity, commonsense knowledge, and an ability to generalize over seen concepts (Costello and Keane, 2000; Wisniewski, 1997). Similar noun compounds might have different meanings based on our common understanding. For example, knowing that “chocolate croissant” means a “croissant filled with chocolate” does not necessarily imply that “chocolate bunny” would mean “a bunny filled with chocolate”, but rather a piece of “chocolate in the shape of a bunny”. Several works have evaluated and analyzed language models on the task of interpreting and predicting the emergence of these noun compounds and found that models generally show a moderate performance (Coil and Shwartz, 2023; Dhar and van der Plas, 2019; Kuznetsova et al., 2013). Other works have successfully trained neural networks to generate neologisms (i.e. newly coined words or phrases) (Das and Ghosh, 2017; Lencione et al., 2022). On the other hand, previous works have also shown that large language models can fail at linguistic generalization tasks such as morphologically deriving new words from nonce roots (Ismayilzada et al., 2024a; Weissweiler et al., 2023) and can occasionally duplicate large amounts of text from its training data (McCoy et al., 2021). Similarly, a recent work explores the linguistic creativity of both large language models and humans by reconstructing their text output from the existing text snippets on the web and finds that the seemingly remarkable creativity of model outputs may be in large part attributable to the remarkable creativity of human-written texts on the web (Lu et al., 2024b).
3.2. Creative Problem-Solving
@@ 3.2.1. Convergent Thinking
Language models have recently been evaluated on problems that require convergent thinking. Lin et al. (2021) tests language models on solving riddles that require creativity and commonsense and finds a significant gap between model and human performance. Naeini et al. (2023) uses the popular British quiz show Only Connect’s Connecting Wall that mimics RAT formulation with built-in, deliberate red herrings (i.e. misleading stimuli or distractors) and evaluates large language models such as GPT-4 on these problems. They report poor model performance and show that models are highly susceptible to distractors in the input and manifest a form of fixation effect (a.k.a functional fixedness or Einstellung effect) (Wiley, 1998; Smith and Blankenship, 1991; Barber, 1960). This type of cognitive bias forces the model to fixate on its past knowledge and prevents it from thinking “out-of-the-box”. The same effect is also found when models are evaluated on everyday problems involving unconventional use of objects (Tian et al., 2023). Very recently, large language models such as GPT-4o have been evaluated on the popular New York Times game Connections and have been found to struggle with associating encyclopedic and linguistic knowledge at an abstract level (Samadarshi et al., 2024). Another study investigating both convergent and divergent creativity of language models has revealed that language models also fall short of demonstrating human-like convergent creativity in code generation (Lu et al., 2024c).
3.2.2. Divergent Thinking
@@
While language models perform strongly on AUT-like divergent thinking tasks, they, however, struggle when these tasks require some form of lateral thinking or “thinking out-of-the-box” (Huang et al., 2024b). For example, recent works have found that defying default commonsense associations and modeling unexpected or unlikely situations are challenging for large language models (Jiang et al., 2023; Tian et al., 2023; Zhao et al., 2023a). Figure 4(a) illustrates a creative problem-solving example from Tian et al. (2023) that involves unconventional use of everyday objects.
3.2.3. Abstraction and Analogy-Making
@@
As the cornerstone of human intelligence, abstraction, and abstract reasoning are typically evaluated using visual IQ tests in humans. Popular examples of these tests are the RAVEN progressive matrices (Raven, 1938), Bongard problems (Bongard, 1970) and the recently introduced Kandinsky Patterns (Holzinger et al., 2019), the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019) and its variations (Moskvichev et al., 2023). These tests require the participants to identify and complete an abstract visual pattern based on given examples. Although several attempts have been made to solve these tasks using both symbolic-based and neural network-driven approaches (Lorello et al., 2024; Mirchandani et al., 2023; Hu et al., 2023; Xu et al., 2022; Santoro et al., 2018), modern AI systems still struggle with solving RAVEN-like (Ahrabian et al., 2024; Gendron et al., 2023; Odouard and Mitchell, 2022; Zhang et al., 2019a) and ARC-like tasks (Moskvichev et al., 2023; Mitchell et al., 2023; Xu et al., 2023; Odouard and Mitchell, 2022; Zhang et al., 2021). Analysis of abstraction via a serial reproduction task (Langlois et al., 2021) where participants are asked to produce a textual stimulus for the next participant upon observing a visual stimulus and vice versa, has suggested that GPT-4 unlike humans relies heavily on linguistic representations even in vision-only paradigm (Kumar et al., 2024). Figure 5 illustrates an example from the ARC task (Chollet, 2019). The problems in this corpus are quite hard to solve to the extent that this task has been recognized as the de facto benchmark for measuring progress towards Artificial General Intelligence (AGI) and a public competition with a grand prize of $ has recently been launched222https://arcprize.org/. At the time of writing this paper, the highest score is far from the passing threshold of (human-level).
Analogy-Making
@@
In its basic form, analogy-making is the ability to identify a relation between two concepts and apply it to a new concept. For example, Paris is to France as Tokyo is to Japan (i.e. capital:country relation). Early approaches to computational analogy-making were symbolic-based and required extensive hand-coded input i.e. structured representations of both the entities and their relations (Turney, 2008; Gentner, 1983; Falkenhainer et al., 1989). Later, word embedding models based on neural networks were shown to exhibit analogy-making abilities at the word level and most works focused on a limited set of analogy types based on a handful of relations that are often of a morphological nature. (Marquer et al., 2022; Gladkova et al., 2016; Mikolov et al., 2013). These do not encompass the typical analogical reasoning humans perform in everyday life about complex situations. More recent work has focused on including a multitude of relations and datasets that are used to test analogical reasoning in humans. (Petersen and van der Plas, 2023; Ushio et al., 2021; Jacob et al., 2023)
While some works have argued for emergent analogical reasoning abilities of large language models (Webb et al., 2022; Yasunaga et al., 2023; Hu et al., 2023), other works have shown that these models lack the robustness and generality exhibited by humans when it comes to long text analogies (Wijesiriwardene et al., 2023; Zhu and de Melo, 2020), scientific analogies (Yuan et al., 2023; Czinczoll et al., 2022), story analogies (Jiayang et al., 2023; Sourati et al., 2023; Sultan and Shahaf, 2022; Nagarajah et al., 2022), visual analogies (Opielka et al., 2024; Bitton et al., 2022; Zhang et al., 2019a) and complex analogical reasoning (Musker et al., 2024; Lewis and Mitchell, 2024). Figure 4(b) illustrates an example from the analogical reasoning over narratives benchmark (Sourati et al., 2023).
3.3. Artistic Creativity
@@ 3.3.1. Story Generation
Language models have also been evaluated on their ability to produce and judge creative content as a professional writer (Chakrabarty et al., 2023b; Marco et al., 2024; Gómez-Rodríguez and Williams, 2023). Chakrabarty et al. (2023b) generates short stories from LLMs based on the plots of popular fictional stories published in the New York Times and conducts a fine-grained expert assessment of both model-generated and original stories. Their study shows that LLMs significantly lag behind seasoned writers in producing inherently creative content. Studies also demonstrate that LLMs are unreliable evaluators of creativity (Chhun et al., 2024; Chakrabarty et al., 2023b). Additionally, (Tian et al., 2024) finds that LLM-generated stories are positively homogenous and typically lack suspense and tension. LLMs have also been shown to produce more complex, but less creative stories than average humans (Ismayilzada et al., 2024b).
@@ 3.4. Scientific Creativity
Another line of work has focused on automating the discovery of other scientific knowledge such as generating new mathematical conjectures or theories (Raayoni et al., 2021; Wu and Tegmark, 2019; Chen et al., 2016; Buchberger et al., 2006; Fajtlowicz, 1988), automatically proving theorems (Hubert et al., 2024b; Trinh et al., 2024), discovering new concepts (Hakuk and Reich, 2020; Iten et al., 2020; Lenat and Brown, 1984) and predicting new molecular structures (Abramson et al., 2024; Zambaldi et al., 2024; Jumper et al., 2021; Lindsay, 1980) among others. A notable example in this area is the recent AlphaFold model (Jumper et al., 2021) that can predict millions of intricate 3D protein structures which has the potential to significantly accelerate research in biology.
@@
While the latest advancements in the automation of scientific creativity are remarkable, these results should be taken with a grain of salt. Most of the recent end-to-end automation frameworks are powered by LLMs, hence, they face the same challenges and issues we discussed in the previous sections such as hallucinations, lack of content diversity, novelty, and robust reasoning capabilities. For example, one of the aforementioned large-scale idea generation studies (Si et al., 2024) finds that out of LLM-generated ideas only are unique. Their qualitative analysis also reveals some common failure modes such as vague implementation details, misuse of datasets, inappropriate baselines, unrealistic assumptions, and overall poorly-motivated ideas. Similarly, another study benchmarking the machine learning experimentation capabilities of LLMs reports hallucinations and poor planning as some of the major issues with these models (Huang et al., 2024c).
@@ 5.1.2. Dimensions of Creativity
As we discussed earlier, there are many dimensions of creativity, but most works generally focus on evaluation of the novelty and usefulness dimensions. However, surprise, agency and spontaneity dimensions are also equally important. Humans typically communicate an emotion or a deeper meaning through creative products and their creative process is characterized by spontaneous “Aha” or “Eureka” moments coupled with deliberate decisions made at each step of the way 666https://www.newyorker.com/culture/the-weekend-essay/why-ai-isnt-going-to-make-art. However, current AI systems lack agency and are typically trained to generate the most likely output leaving no room for any intentional or spontaneous action (Peeperkorn et al., 2023; Franceschelli and Musolesi, 2023). Therefore, a holistic evaluation of machine creativity should involve consideration of all these different dimensions that characterize human creativity.
@@ Acknowledgements.
We gratefully acknowledge the support of Swiss National Science Foundation (grant 205121_207437: C - LING). We also thank Angelika Romanou for her help with creating figures for this paper.References
@@
- Arnheim (2001) Rudolf Arnheim. 2001. What it means to be creative. British Journal of Aesthetics 41 (2001), 24–25. https://api.semanticscholar.org/CorpusID:191484364
@@
- Barber (1960) Theodore Xenophon Barber. 1960. Rigidity of Behavior. A Variational Approach to the Effect of Einstellung.Abraham S. Luchins , Edith Hirsch Luchins. The Quarterly Review of Biology 35 (1960), 255–255. https://api.semanticscholar.org/CorpusID:85022531
- Barron (1955) Frank Barron. 1955. The disposition toward originality. Journal of abnormal psychology 51 3 (1955), 478–85. https://api.semanticscholar.org/CorpusID:30102436
@@
- Elgammal et al. (2017) A. Elgammal, Bingchen Liu, Mohamed Elhoseiny, and Marian Mazzone. 2017. CAN: Creative Adversarial Networks, Generating ”Art” by Learning About Styles and Deviating from Style Norms. In International Conference on Innovative Computing and Cloud Computing. https://api.semanticscholar.org/CorpusID:24986117
@@
- Gómez-Rodríguez and Williams (2023) Carlos Gómez-Rodríguez and Paul Williams. 2023. A Confederacy of Models: a Comprehensive Evaluation of LLMs on Creative Writing. In Findings of the Association for Computational Linguistics: EMNLP 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 14504–14528. https://doi.org/10.18653/v1/2023.findings-emnlp.966
@@
- Huang et al. (2024c) Qian Huang, Jian Vora, Percy Liang, and Jure Leskovec. 2024c. MLAgentBench: Evaluating Language Agents on Machine Learning Experimentation. arXiv:2310.03302 [cs.LG] https://arxiv.org/abs/2310.03302
- Huang et al. (2024b) Shulin Huang, Shirong Ma, Yinghui Li, Mengzuo Huang, Wuhe Zou, Weidong Zhang, and Haitao Zheng. 2024b. LatEval: An Interactive LLMs Evaluation Benchmark with Incomplete Information from Lateral Thinking Puzzles. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), Nicoletta Calzolari, Min-Yen Kan, Veronique Hoste, Alessandro Lenci, Sakriani Sakti, and Nianwen Xue (Eds.). ELRA and ICCL, Torino, Italia, 10186–10197. https://aclanthology.org/2024.lrec-main.889
@@
- Ismayilzada et al. (2024a) Mete Ismayilzada, Defne Circi, Jonne Sälevä, Hale Sirin, Abdullatif Köksal, Bhuwan Dhingra, Antoine Bosselut, Lonneke van der Plas, and Duygu Ataman. 2024a. Evaluating Morphological Compositional Generalization in Large Language Models. arXiv:2410.12656 [cs.CL] https://arxiv.org/abs/2410.12656
@@
- Ismayilzada et al. (2024b) Mete Ismayilzada, Claire Stevenson, and Lonneke van der Plas. 2024b. Evaluating Creative Short Story Generation in Humans and Large Language Models. arXiv:2411.02316 [cs.CL] https://arxiv.org/abs/2411.02316
@@
- Peeperkorn et al. (2023) Max Peeperkorn, Dan Brown, and Anna Jordanous. 2023. On characterizations of large language models and creativity evaluation. (2023).
@@
- Smith and Blankenship (1991) Steven M. Smith and Steven E. Blankenship. 1991. Incubation and the persistence of fixation in problem solving. The American journal of psychology 104 1 (1991), 61–87. https://api.semanticscholar.org/CorpusID:10359632
@@
- Stein (1953) Morris Isaac Stein. 1953. Creativity and Culture. Creativity in Art, Religion, and Culture (1953). https://api.semanticscholar.org/CorpusID:144504562
@@
- Weisberg (2015a) Robert W. Weisberg. 2015a. On the Usefulness of “Value” in the Definition of Creativity. Creativity Research Journal 27, 2 (2015), 111–124. https://doi.org/10.1080/10400419.2015.1030320 arXiv:https://doi.org/10.1080/10400419.2015.1030320
- Weisberg (2015b) Robert W. Weisberg. 2015b. On the Usefulness of “Value” in the Definition of Creativity. Creativity Research Journal 27 (2015), 111 – 124. https://api.semanticscholar.org/CorpusID:146831140
@@
- Wiley (1998) Jennifer Wiley. 1998. Expertise as mental set: The effects of domain knowledge in creative problem solving. Memory & Cognition 26 (1998), 716–730. https://api.semanticscholar.org/CorpusID:45775509
@@
- Xie et al. (2023c) Binzhu Xie, Sicheng Zhang, Zitang Zhou, Bo Li, Yuanhan Zhang, Jack Hessel, Jingkang Yang, and Ziwei Liu. 2023c. FunQA: Towards Surprising Video Comprehension. ArXiv abs/2306.14899 (2023). https://api.semanticscholar.org/CorpusID:259262297
@@
- Xie et al. (2023a) Zhuohan Xie, Trevor Cohn, and Jey Han Lau. 2023a. The Next Chapter: A Study of Large Language Models in Storytelling. In Proceedings of the 16th International Natural Language Generation Conference, C. Maria Keet, Hung-Yi Lee, and Sina Zarrieß (Eds.). Association for Computational Linguistics, Prague, Czechia, 323–351. https://doi.org/10.18653/v1/2023.inlg-main.23
- Xie et al. (2023b) Zhuohan Xie, Miao Li, Trevor Cohn, and Jey Lau. 2023b. DeltaScore: Fine-Grained Story Evaluation with Perturbations. In Findings of the Association for Computational Linguistics: EMNLP 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 5317–5331. https://doi.org/10.18653/v1/2023.findings-emnlp.353
@@
- Yang and Jin (2024) Dingyi Yang and Qin Jin. 2024. What makes a good story and how can we measure it? a comprehensive survey of story evaluation. arXiv preprint arXiv:2408.14622 (2024).